# Cryogenic Microwave Whispering Gallery Mode Spectroscopy of Paramagnetic Impurities in High-Purity Crystalline LiF

### v1

@@ Abstract
A low-noise cryogenic microwave spectroscopy experiment was performed on a high-purity lithium fluoride (LiF) crystal. The spectroscopy data revealed avoided level crossing interactions in whispering gallery modes, indicative of electron spin resonance (ESR) coupling with paramagnetic impurities. Analysis of the interaction spectra identified distinct spin systems corresponding to ${S=3/2,I=7/2}$, ${S=1,I=7/2}$, and ${S=3/2,I=0}$. The number of hyperfine splittings observed, together with the natural abundance of ions possessing the appropriate nuclear spin values, suggest that V2+ and V3+ impurities, exhibiting orthorhombic distortion, are the most likely sources of the narrow interaction features. This interpretation is supported by earlier ESR studies Chan and Shields (1970) and established manufacturing records for LiF crystal growth Anbinder, Zakhariya, and Zhikhareva (1973). Additionally, a separate set of broader interaction points is consistent with an orthorhombic model involving a ${S=3/2,I=0}$ spin system, although the specific impurity responsible for this interaction remains unidentified.
@@ II Experiment
The LiF cylinder was grown epitaxially in the z direction then cut along the (001) plane. It has a diameter of 30 mm and a thickness of 30 mm. In addition, a 3 mm by 15 mm hole is precisely drilled into the bottom of the crystal to accommodate a copper rod from the cavity. Looped conductors are attached on opposite sides of the copper cavity and are responsible for coupling RF signals into and out of the cavity. The position of the probes are adjusted such that reflected power from the cavity due to impedence mismatch is minimized. The cylindrical cavity itself is manufactured from oxygen-free high thermal conductivity copper. This is to attain high quality factor WGMs at low temperatures by removing hot spots from oxygen impurities that increase surface losses.
@@ III Theoretical Model of the Experiment
Here, $f_{WG}$ is the frequency of the WG mode and $\Delta E$ is the energy difference between the $m$th and $n$th eigenstates of (1). For specific impurity populations, an orthorhombic spin model can be considered. In this model, it is sufficient to consider only the electron Zeeman interaction, zero-field splitting interaction, and the hyperfine interaction as the superhyperfine interaction is negligible Chan, Shields, and Weston (1971):
Figure 2: Simulated EPR spectra generated using equation (III) with parameter values for $g$, $E$, $D$, and $A$ taken from Table 1. The spectra are fitted to selected interaction points and are categorized by their spectroscopic signal-to-noise ratio (SNR). Simulations assume different electron and nuclear spin numbers for distinct sets of data points. Each spectrum is assigned either a most likely impurity type or labeled as unknown (denoted by X).
$\hat{H}$ $=\mu_{B}g_{e}B_{z}\hat{S}_{z}+hD\left(S_{z}^{2}-\frac{1}{3}S(S+1)\right)$
$+hE(S_{x}^{2}-S_{y}^{2})+A\mathbf{I}\cdot\mathbf{S}$ (3)
Here, the first term is the electron Zeeman interaction, the 2nd and 3rd terms are the zero-field splitting interactions and the 4th term is the hyperfine interaction. The Hamiltonian in (III) is termed the orthorhombic spin Hamiltonian and is used to model orthorhombic distortions on crystal lattice sites that are caused by ion impurities such as V2+ Chan and Shields (1968).
@@ IV Main Results and Discussions
F-F-F-V2+F-F-F-$\mathbf{B}$$\mathbf{x}$$\mathbf{z}$$\mathbf{y}$
Figure 4: V2+ orthorhombic distortion in a LiF crystal lattice. The surrounding fluoride ions shift slightly along the crystal axes, creating three distinct distances to the central V2+ ion, rather than a uniform distance as in a cubic field.
During certain manufacturing processes of pure LiF, impurities such as vanadium and cobalt, both possessing a nuclear spin of $I=7/2$, are unavoidably introduced into the sample at low concentrations Anbinder, Zakhariya, and Zhikhareva (1973). The most general form of vanadium crystal field distortion in LiF is orthorhombic Chan and Shields (1970). Therefore, the Hamiltonian in equation (III) was used to simulate the spin system. This Hamiltonian was diagonalized to obtain the energy levels and corresponding EPR spectra, which are represented by the dashed lines in Figure 2. Forbidden transitions can be filtered out using equation (5). The parameters $g$, $E$, $D$, and $A$ were treated as free parameters to fit the simulated spectra to the interaction points observed in the spectroscopy. The best-fit values, along with those reported in Chan, Shields, and Weston (1971), are listed in Table 1.
Note that the parameter values determined from this ESR experiment differ significantly from those reported in Chan, Shields, and Weston (1971). Although large deviations in these parameters for the same host and impurity system are uncommon, fundamental differences in the defect environment and crystal geometry relative to the applied magnetic field in ESR spectroscopy may account for such discrepancies. For example, different charge compensation mechanisms, such as the substitution of V2+ for Li+, which introduces a net $+1$ charge, may be compensated by F- vacancies, Li+ interstitials, or nearby impurity ions such as OH- or O2- Watkins (1959). These mechanisms can distort the crystal lattice in different ways, altering the local symmetry—which affects the $g$-tensor and zero-field splitting parameters, and/or the covalency, which influences the hyperfine coupling constant. It is also possible that the LiF sample used in this study consists of multiple domains with slightly different orientations due to non-ideal fabrication conditions, such as uneven cooling or mechanical strain, a phenomenon known as mosaicity Poole (1967).
@@
$E$ (GHz) $1.9\pm 0.1$ (±)plus-or-minus($\pm$)( ± )4.80
$D$ (GHz) $3.1\pm 0.1$ ()minus-or-plus($\mp$)( ∓ )1.04
$A$ (GHz) $2.7\pm 0.01$ (±)plus-or-minus($\pm$)( ± )0.25 $\pm$ 0.005

### v2

@@ Abstract
A low-noise cryogenic microwave spectroscopy experiment was performed on a high-purity lithium fluoride (LiF) crystal. The spectroscopy data revealed avoided level crossing interactions in whispering gallery modes, indicative of electron spin resonance (ESR) coupling with paramagnetic impurities. Analysis of the interaction spectra identified distinct spin systems corresponding to $(S=3/2,I=7/2)$, $(S=1,I=7/2)$, and $(S=3/2,I=0)$. The number of hyperfine splittings observed, together with the natural abundance of ions possessing the appropriate nuclear spin values, suggest that V2+ and V3+ impurities, exhibiting orthorhombic distortion, are the most likely sources of the narrow interaction features. This interpretation is supported by earlier ESR studies and established manufacturing records for LiF crystal growth. Additionally, a separate set of broader interaction points is consistent with an orthorhombic model involving a $(S=3/2,I=0)$ spin system, although the specific impurity responsible for this interaction remains unidentified.
@@ II Experiment
The LiF cylinder was grown epitaxially in the z direction using the Czochralski technique then cut along the (001) plane. It has a diameter of 30 mm and a thickness of 30 mm. In addition, a 3 mm by 15 mm hole is precisely drilled into the bottom of the crystal to accommodate a copper rod from the cavity. Looped conductors are attached on opposite sides of the copper cavity and are responsible for coupling RF signals into and out of the cavity. The position of the probes are adjusted such that reflected power from the cavity due to impedence mismatch is minimized. The cylindrical cavity itself is manufactured from oxygen-free high thermal conductivity copper. This is to attain high quality factor WGMs at low temperatures by removing hot spots from oxygen impurities that increase surface losses.
@@ III Theoretical Model of the Experiment
Here, $f_{WG}$ is the frequency of the WG mode and $\Delta E$ is the energy difference between the $m$th and $n$th eigenstates of (1). For specific impurity populations, an orthorhombic spin model can be considered. In this model, it is sufficient to consider only the electron Zeeman interaction, zero-field splitting interaction, and the hyperfine interaction as the super-hyperfine interaction is negligible Chan, Shields, and Weston (1971). For arbitrary orientations of orthorhombic distortion, the spin interaction Hamiltonian can be written as follows Nehrkorn et al. (2015):
$\hat{H}$ $=\mu_{B}g_{e}B_{z}\hat{S}_{z}+h\mathbf{\hat{S}}^{T}\cdot\mathbf{D}\cdot\mathbf{\hat{S}}+hA\mathbf{I}\cdot\mathbf{S}$ (3)
Figure 2: Simulated EPR spectra generated using equation (3) with parameter values for $g$, $E$, $D$, and $A$ taken from Table 1. The spectra are fitted to selected interaction points and are categorized by their spectroscopic signal-to-noise ratio (SNR). Simulations assume different electron and nuclear spin numbers for distinct sets of data points. Each spectrum is assigned either a most likely impurity type or labeled as unknown (denoted by X).
Here, the first term represents the electron Zeeman interaction, the second term corresponds to the zero-field splitting interaction, and the third term describes the hyperfine interaction. For orthorhombic distortions, $\mathbf{D}$ can be oriented to the principal axis to reduce the zero-field splitting term in Equation (3) to the following Hutton (1969); König and Schnakig (1976):
$\mathbf{\hat{S}}^{T}\cdot\mathbf{D}\cdot\mathbf{\hat{S}}$ $=D\left(S_{z}^{2}-\frac{1}{3}S(S+1)\right)+E(S_{x}^{2}-S_{y}^{2})$ (4)
@@ IV Main Results and Discussions
F-F-F-V2+F-F-F-$\mathbf{B}$$\mathbf{x}$$\mathbf{z}$$\mathbf{y}$$\mathbf{D}_{zz}$$\mathbf{D}_{yy}$$\mathbf{D}_{xx}$
Figure 4: Orthorhombic distortion of a V2+ ion in the LiF crystal lattice. The surrounding fluoride ions are slightly displaced along the crystal axes, resulting in three distinct distances to the central V2+ ion, as opposed to the uniform distances found in a cubic field. The orientation of the zero-field tensor in its principal axis frame is also shown (right) for all impurities listed in Table 1. Each axis of the tensor is labeled as $\mathbf{D}_{xx}$, $\mathbf{D}_{yy}$, or $\mathbf{D}_{zz}$.
During Czochralski crystal growth, trace impurities such as vanadium and cobalt—both possessing a nuclear spin of $I=7/2$—can be unintentionally incorporated into the sample at low concentrations. A well-documented case of this occurs in the growth of GaAs, where vanadium originating from a boron nitride crucible was found to contaminate the crystal Brandt et al. (1985). The most general form of vanadium crystal field distortion in LiF is orthorhombic Chan and Shields (1970). Therefore, the Hamiltonian in equation (3) was used to simulate the spin system. This Hamiltonian was diagonalized to obtain the energy levels and corresponding EPR spectra, which are represented by the dashed lines in Figure 2. Forbidden transitions can be filtered out using equation (6). The parameters $g$, $E$, $D$, and $A$ were treated as free variables to fit the simulated spectra to the interaction points observed in the experimental data. The resulting $\mathbf{D}$ tensor was then rotated into its principal axis frame (see Figure 4 for the orientation of the $\mathbf{D}$ tensor for all impurities relative to the laboratory frame). The best-fit values, along with those reported in Chan, Shields, and Weston (1971), are listed in Table 1.
Note that some of the parameter values obtained in this ESR experiment differ slightly from those reported in Chan, Shields, and Weston (1971). These discrepancies may arise from fundamental differences in the defect environment and crystal geometry relative to the applied magnetic field in ESR spectroscopy. For example, different charge compensation mechanisms—such as the substitution of V2+ for Li+, which introduces a net $+1$ charge—may be compensated by F- vacancies, Li+ interstitials, or nearby impurity ions such as OH- or O2- Watkins (1959). These mechanisms can distort the crystal lattice in various ways, modifying the local symmetry and thereby affecting the $g$-tensor and zero-field splitting parameters, and/or altering the covalency, which in turn influences the hyperfine coupling constant.
@@
$E$ (GHz) (±$(\pm)0.58\pm 0.1$ (±)plus-or-minus($\pm$)( ± )0.348
$D$ (GHz) $(\mp)4.5\pm 0.1$ ()minus-or-plus($\mp$)( ∓ )4.80
$A$ (GHz) $0.27\pm 0.01$ (±)plus-or-minus($\pm$)( ± )0.25 $\pm$ 0.005
@@ IV.1 Coupling Strength and Spin Q Factor
The following expression is used to estimate the concentration of defect spins $n$ from the coupling strength $g_{CS}$, assuming uniform spin–mode coupling and a single dominant impurity species Bushev et al. (2011):
$g_{CS}=g\mu_{B}\sqrt{\frac{\mu_{0}\omega_{0}n\xi_{\perp}}{4\hbar}}$ (9)
@@
Here, $\mu_{B}$ is the Bohr magneton, $\mu_{0}$ is the permeability of free space, $\omega_{0}$ is the angular frequency of the microwave mode, $g$ is the effective Landé $g$-factor, and $\xi_{\perp}$ is the perpendicular magnetic filling factor ($\xi_{\perp}\approx 1$ for TM-WGMs) Bourhill (2016). Since the high SNR narrow interactions are attributed to V2+ spin impurities, the corresponding spin concentration is estimated from equation (9) to be $n=(1.0\pm 0.2)\times 10^{10}$ cm-3. To assess how this compares to the lowest concentration detectable by the current setup, following parameter values are assumed: $g_{CS}=5$ Hz, $\omega_{0}=2\pi\times 20$ GHz, and $\xi=1$, which yields an approximate minimum detectable concentration of $n_{\text{min}}\approx 10^{7}$ cm-3.
@@ V Conclusion
Author Declarations
This work was funded by the ARC Centre of Excellence for Engineered Quantum Systems, CE170100009, and Dark Matter Particle Physics, CE200100008.
Conflict of Interest
The authors have no conflicts to disclose.
Author Contributions
Steven Samuels: Writing - original draft (lead); Writing - review & editing (equal). William Campbell: Conceptualization (equal); Writing - review & editing (equal). Michael E. Tobar: Conceptualization (equal); Writing - review & editing (equal). Maxim Goryachev: Conceptualization (equal); Writing - review & editing (equal).
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.